OpenStudy (samsan9):

which conic section does the equation below describe? (x+1)^2+(y-3)^2=4 A.)Hyperbola B.)Circle C.)Parabola D.)Ellipse

OpenStudy (bradely):

(x+1)^2+(y-3)^2=4 Simplified form (x-(-1))^2+(y-3)^2=2^2 The equation is in the form (x-h)^2+(y-h)^2=r^2 So conic is circle B is correct Source:www.mathskey.com

OpenStudy (kainui):

It looks like the pythagorean theorem. But what's going on with it? The hypotenuse is constant, and the sides can change. Can you try to imagine picking a point, then extending a constant length from it and then imagining there being a bunch of different possible triangles that you can make as you move the hypotenuse to different places? 

OpenStudy (bradely):

hi, any pre-calculus help please post on source site